OpenStudy (anonymous):

how do you solve this -3 √ -1000

OpenStudy (anonymous):

is that the negative third root of -1000?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

=(-1000)^(-1/3) = 1/((-1000)^(1/3)=1/-10 = -1/10

OpenStudy (anonymous):

because the cube root of a negative is negative and the cube root of 1000 is 10

OpenStudy (anonymous):

I replaced the cube root with the fractional exponent, is that okay by you?

OpenStudy (anonymous):

yes so the -3 = -1 and the -1000=-10 how did you get the positive

OpenStudy (anonymous):

When you move a factor from the numerator to the denominator, the sign on the exponent switches. So A^-(1/3) = 1/A^(1/3)

OpenStudy (anonymous):

oic ty

OpenStudy (anonymous):

so the answer is -1 √ 10 or 1 √ -10